OpenStudy (anonymous):

Calculate dy/dx.You need not expand your answer. y = x/(x − 7)(x − 4)

myininaya (myininaya):

not sure what all is in the denominator

OpenStudy (anonymous):

\[\frac{x}{(x-7)(x-4)} ?\]

myininaya (myininaya):

thats what i kindof thought he meant

myininaya (myininaya):

use quotient rule is one way

OpenStudy (anonymous):

Its dealing with derivatives I'm not sure what to use and how to solve it.

myininaya (myininaya):

you have f/g so use quotient rule (f'g-g'f)/(g^2)

myininaya (myininaya):

I will show you three ways! ---------------------- 1st way: Quotient Rule! \[y=\frac{x}{(x-7)(x-4)}=\frac{x}{x^2-11x+28}\] \[y'=\frac{(x)'(x^2-11x+28)-x(x^2-11x+28)'}{(x^2-11x+28)^2}\] \[=\frac{1(x^2-11x+28)-x(2x-11+0)}{(x^2-11x+28)^2}\] ------------------------ 2nd way: Product Rule! \[y=x(x^2-11x+28)^{-1}\] \[y'=(x)'(x^2-11x+28)^{-1}+x[(x^2-11x+28)^{-1}]'\] \[y'=1(x^2-11x+28)^{-1}+x(-1)(x^2-11x+28)^{-1-1}(x^2-11x+28)'\] \[=\frac{1}{x^2-11x+28}-x \frac{1}{(x^2-11x+28)^2}(2x-11+0)\] ------------------------ 3rd way: Logarithmic Differentiation (my favorite!) \[y=\frac{x}{(x-7)(x-4)}\] Before Differentiating both sides we will do ln( ) of both sides \[\ln(y)=\ln(\frac{x}{(x-7)(x-4)})\] \[\ln(y)=\ln(x)-\ln[(x-7)(x-4)]\] \[\ln(y)=\ln(x)-[\ln(x-7)+\ln(x-4)]\] \[\ln(y)=\ln(x)-\ln(x-7)-\ln(x-4)\] Now the derivative part! \[\frac{y'}{y}=\frac{1}{x}-\frac{1}{x-7}-\frac{1}{x-4}\] Now we wanted y' so to isolate y' multiply y on both sides \[y'=y(\frac{1}{x}-\frac{1}{x-7}-\frac{1}{x-4})\] But remember \[y=\frac{x}{(x-7)(x-4)}\] \[y'=\frac{x}{(x-7)(x-4)} (\frac{1}{x}-\frac{1}{x-7}-\frac{1}{x-4})\]